OpenStudy (shelby1290):

The height of a soccer ball is modelled by h(t) = -4.9t^2+ 19.6t + 0.5, where height h(t) is in metres and time, t, in seconds. a) What is the maximum height of the ball?

OpenStudy (shelby1290):

h(t) = -4.9t^2 + 19.6t + 0.5 = -4.9(t^2-4t)+0.5 -4 divided by 2 = 2^2 = 4 =-4.9(t^2-4t+4-4)+0.5 =-4.9(t^2-4t+4-4)+0.5 = -4.9(t-4t+4)+0.5+19.6 =-4.9(t-2)^2+20.1

OpenStudy (shelby1290):

The maximum height is 20.1 m

OpenStudy (shelby1290):

Is my answer correct?

Mehek (mehek14):

yup correct

OpenStudy (shelby1290):

Can I get the same answer with the formula that you posted before?

OpenStudy (shelby1290):

I think it was.... -2 / ab

Mehek (mehek14):

another way you could've done was \(\tt{\dfrac{-b}{2a}}\) \(\dfrac{-19.6}{2*(-4.9)}\\ \dfrac{-19.6}{-9.8} = 2 \) \(h(2) = -4.9*2^2+ 19.6*2 + 0.5\\ -4.9*4+39.2\\-19.6 + 39.2 + 0.5 = 20.1\)

Mehek (mehek14):

yes both ways work and its \(\tt{\dfrac{-b}{2a}}\)

OpenStudy (shelby1290):

Thanks for the help!

Mehek (mehek14):

anytime \(\ddot\smile\)